OpenStudy (anonymous):

The functions f(x) and g(x) are described below: f(x) = 7x + 9 g(x) = 7x - 4 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units.

OpenStudy (anonymous):

13 units. 9 - (-4) = 9 + 4 =13

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (triciaal):